OpenStudy (anonymous):

What is the equation of the line passing through the point (-2,-1) and perpendicular to the line y=1/3x+2?

OpenStudy (anonymous):

y = 1/3x + 2. You are looking for the perpendicular line to this. That means the slope is going to be the negative reciprocal of the slope in the given equation. Do you know what the slope is in y = 1/3x + 2 ?

OpenStudy (anonymous):

1/3?

OpenStudy (anonymous):

correct. But we are looking for the perpendicular line, which means we need the negative reciprocal of the slope 1/3. What that means is we " flip " the slope and change the sign. Which means 1/3 slope for a perpendicular line becomes -3/1 or just -3. now we will use point slope formula because we have the point and the slope. y - y1 = m(x - x1) using slope(m) = -3 (-2,-1) x1 = -2 and y1 = -1 now just plug your info into the formula and you will have your answer.